# Resonant Type-II + Type-I Hybrid Leptogenesis in SO(10) with Triplet–Neutrino Mass Degeneracy $M_{T}\simeq M_{N_{3}}$

### v2

@@ V Numerical Results
Figure 5: Hybrid resonant leptogenesis in the plane of the hybrid CP phase $\phi_{\rm HR}$ and the heavy scale $\Lambda\sim M_{T}\sim M_{N_{3}}$. Left panel: colour shading shows a proxy for the baryon asymmetry $Y_{B}$ generated through the resonant interference of type-I and type-II contributions; the horizontal band indicates the region compatible with the observed baryon asymmetry. Right panel: indirect collider sensitivity to the same hybrid CP phase through CP-odd Higgs observables, illustrating the link between high-scale baryogenesis and precision measurements at the LHC.
Figure 5 provides a unified view of the hybrid leptogenesis mechanism and its experimental implications. The left panel displays the generation of the baryon asymmetry in the $\phi_{\rm HR}$$\Lambda$ plane, showing that successful leptogenesis occurs naturally for quasi-degenerate triplet and right-handed neutrino masses in the presence of a non-vanishing hybrid CP phase. The horizontal band highlights the region compatible with the observed baryon asymmetry, demonstrating that the hybrid resonant mechanism efficiently reproduces the correct order of magnitude without extreme parameter tuning.
The right panel shows the indirect collider sensitivity to the same hybrid CP phase through CP-odd Higgs observables induced by dimension-six operators after integrating out the heavy states. Although the leptogenesis scale lies far above the collider energy, the figure illustrates that sizable CP-violating effects persist over a wide range of the parameter space and can be probed through precision Higgs measurements at the LHC and its high-luminosity upgrade. Taken together, the two panels establish a direct conceptual link between the origin of the baryon asymmetry and experimentally accessible CP-violating observables.
Figure 6: Hybrid resonant leptogenesis parameter space in the plane of the hybrid CP phase $\phi_{\rm HR}$ and the degeneracy parameter $|\delta|=|M_{T}-M_{N_{3}}|/M_{N_{3}}$. The red island indicates the region where the resonant interference between type-I and type-II contributions generates a baryon asymmetry of the observed order. The hatched region is excluded by electric dipole moment constraints, while the blue contour denotes the approximate sensitivity of indirect collider CP-odd observables. The surviving overlap highlights the simultaneous viability and testability of the hybrid leptogenesis framework.
Figure 6 summarizes the phenomenological implications of the hybrid leptogenesis scenario in a single parameter-space scan. The finite red island arises from the combined requirements of resonant enhancement, quasi-degeneracy of the heavy states, and a non-vanishing hybrid CP phase, leading to a predictive region compatible with the observed baryon asymmetry. Electric dipole moment constraints exclude part of the parameter space but do not eliminate the viable region, while indirect collider sensitivity through CP-odd Higgs observables overlaps significantly with the baryogenesis-favoured domain. This demonstrates that hybrid resonant leptogenesis remains both viable and experimentally testable despite originating at very high scales.
Figure 7: Scaling of the baryon asymmetry with the degeneracy parameter $|\delta|$. Dots represent numerical scan points, while the solid line shows a power-law fit, illustrating the resonant enhancement as the heavy states approach degeneracy.
Figure 8: Linear correlation between the CP-odd Higgs observable and $\sin\phi_{\rm HR}$. The solid line represents a linear fit to the numerical points, demonstrating that the same hybrid CP phase governing leptogenesis induces approximately linear effects in collider observables.
Figure 9: Renormalization-group evolution of the effective CP-violating operator coefficient. Dots denote the evolved values at different scales, while the solid line shows a logarithmic fit, indicating mild RG enhancement of low-energy CP-odd observables.
Figure 7 shows the dependence of the baryon asymmetry on the degeneracy parameter $|\delta|=|M_{T}-M_{N_{3}}|/M_{N_{3}}$ obtained from a numerical scan. The discrete points correspond to individual parameter choices, while the solid line represents a power-law fit. The approximate scaling behaviour reflects the resonant enhancement of the CP asymmetry as the heavy states become increasingly degenerate, providing a quantitative characterization of the resonance structure underlying hybrid leptogenesis. Figure 8 illustrates the correlation between the CP-odd Higgs observable and $\sin\phi_{\rm HR}$. The numerical points are well described by a linear fit, indicating that the dominant contribution to the observable arises from the hybrid CP-violating phase. This behaviour supports the interpretation that the same CP phase responsible for baryogenesis induces approximately linear CP-violating effects in Higgs processes accessible at the LHC. Figure 9 displays the renormalization-group evolution of the effective CP-violating operator coefficient between the leptogenesis scale and the electroweak scale. The fitted logarithmic curve captures the leading RG behaviour and shows that the running induces only a mild enhancement of the coefficient. This justifies the use of leading-logarithmic RG evolution in estimating low-energy CP-odd observables while preserving their dependence on the hybrid CP phase.
@@ VI Conclusions
We have developed the first complete formulation of resonant hybrid type-I plus type-II leptogenesis within a renormalizable SO(10) grand unified theory featuring a controlled quasi-degeneracy between the scalar triplet and the heaviest right-handed neutrino, $M_{T}\simeq M_{N_{3}}$. In this regime, the interference between type-I and type-II decay amplitudes induces a genuinely new CP-violating invariant, $\phi_{\rm HR}=\arg(\mu fY_{\nu}^{\dagger})$, which cannot be removed by field redefinitions and governs the dynamics of the baryon asymmetry.
We have shown that this hybrid CP phase leads to a resonant enhancement of the lepton asymmetry through the constructive interference of triplet and right-handed neutrino contributions. As a result, the observed baryon asymmetry of the Universe can be reproduced for mass scales around $10^{11}\,\mathrm{GeV}$ with moderate Yukawa couplings and without invoking large hierarchies or finely tuned mass splittings. The same interactions that generate the baryon asymmetry also give rise to correlated low-energy signatures, including charged-lepton flavour violation and leptonic electric dipole moments, rendering the framework experimentally testable.
The resonant hybrid regime identified here is theoretically well motivated by the structure of SO(10) unification, phenomenologically rich, and highly predictive. It therefore provides a compelling and falsifiable pathway to baryogenesis, linking the origin of the baryon asymmetry to neutrino masses and low-energy CP violation within a unified framework. Finally, we have shown that although the heavy states responsible for hybrid leptogenesis reside far above the kinematic reach of the LHC, their effects do not completely decouple from collider observables. The hybrid CP-violating invariant induces a characteristic pattern of dimension-six operators, leading to correlated signatures in Higgs CP observables, Higgs–lepton couplings, lepton-flavour-violating dilepton production, and electroweak precision measurements. As illustrated in Fig. 3, these observables exhibit different decoupling behaviours with the heavy scale, highlighting the complementarity of Higgs, flavour, and precision probes in testing the hybrid leptogenesis framework. Figure 5 highlights how the same hybrid CP-violating invariant that generates the baryon asymmetry also controls experimentally accessible CP-odd Higgs observables, thereby establishing a direct link between high-scale leptogenesis and precision collider tests. In particular, CP-odd observables in $h\to\tau^{+}\tau^{-}$ provide the most sensitive indirect collider window on the hybrid CP phase, establishing a direct conceptual link between the origin of the baryon asymmetry and precision measurements at the HL-LHC.
The shaped parameter-space island in Fig. 6 shows that the hybrid CP phase responsible for the baryon asymmetry simultaneously governs EDM constraints and collider-accessible CP-odd observables, rendering the mechanism predictive and testable. The fitted scaling of the baryon asymmetry with the degeneracy parameter quantifies the resonant enhancement underlying the hybrid leptogenesis mechanism.
The approximately linear correlation between CP-odd Higgs observables and the hybrid CP phase demonstrates that collider measurements provide a direct probe of the CP violation responsible for baryogenesis.
The mild logarithmic renormalization-group evolution of the effective CP-violating operators shows that the qualitative link between high-scale leptogenesis and low-energy observables remains robust under running effects.
@@ Acknowledgments
The author thanks the theoretical physics community for ongoing inspiration and resources enabling this research.

### v3

@@ V Numerical Results
Figure 5: Hybrid resonant leptogenesis in the plane of the hybrid CP phase $\phi_{\rm HR}$ and the heavy scale $\Lambda\sim M_{T}\sim M_{N_{3}}$. Left: colour shading indicates the baryon asymmetry with the horizontal band showing the observed range. Right: indirect collider sensitivity through CP-odd Higgs observables.
Figure 5 provides a unified view of the generation of the baryon asymmetry and its indirect experimental implications. Successful leptogenesis occurs for quasi-degenerate heavy states in the presence of a non-vanishing hybrid CP phase, while CP-odd Higgs observables probe overlapping regions of parameter space.
Figure 6: Parameter space in the plane of $\phi_{\rm HR}$ and the degeneracy parameter $|\delta|$. The coloured region corresponds to successful baryogenesis, the hatched region is excluded by EDM constraints, and the contour indicates indirect collider sensitivity.
Figure 6 summarizes the combined impact of baryogenesis, electric dipole moment constraints, and collider sensitivity. The finite allowed region arises from the resonant enhancement and the requirement of a non-zero hybrid CP phase.
Figure 7: Scaling of the baryon asymmetry with the degeneracy parameter $|\delta|$.
Figure 8: Correlation between the CP-odd Higgs observable and $\sin\phi_{\rm HR}$.
Figure 9: Renormalization-group evolution of the effective CP-violating operator.
Figures 79 illustrate the resonant scaling of the baryon asymmetry, the linear dependence of collider CP observables on the hybrid CP phase, and the mild renormalization-group enhancement of low-energy operators. These results quantitatively support the resonant origin of the asymmetry and its correlated phenomenological signatures.
@@ VI Conclusions
We have investigated a regime of hybrid resonant Type-I and Type-II leptogenesis within a renormalizable SO(10) grand unified framework in which the scalar electroweak triplet and the heaviest right-handed neutrino possess comparable masses, $M_{T}\simeq M_{N_{3}}$. Such a quasi-degenerate spectrum can naturally arise in unified constructions where both states originate from common scalar multiplets, and does not rely on ad hoc assumptions. In this regime, the interference between Type-I and Type-II decay amplitudes induces a physical CP-violating invariant, $\phi_{\rm HR}=\arg(\mu fY_{\nu}^{\dagger})$, which governs the generation of the lepton asymmetry and is absent in purely Type-I or Type-II scenarios.
We have shown that this hybrid CP phase leads to a resonant enhancement of the lepton asymmetry through the interplay of triplet- and neutrino-mediated decay channels. For representative benchmark parameters, including explicit treatment of washout effects from lighter right-handed neutrinos, the observed baryon asymmetry of the Universe can be reproduced for mass scales around $10^{11}\,\mathrm{GeV}$ with moderate Yukawa couplings and without requiring extreme mass hierarchies or fine tuning.
The framework consistently incorporates gauge-invariant scalar interactions, flavour effects, and thermal washout dynamics, allowing for a quantitative and reproducible numerical analysis. The resulting baryogenesis mechanism is thus both theoretically well motivated and phenomenologically viable within renormalizable SO(10) models.
Beyond the generation of the baryon asymmetry, the hybrid CP-violating structure gives rise to correlated low-energy signatures. The same invariant $\phi_{\rm HR}$ contributes to charged-lepton flavour violating processes and leptonic electric dipole moments, while also inducing higher-dimensional operators affecting Higgs CP observables and electroweak precision quantities. Although the heavy states responsible for leptogenesis lie far above collider energies, their indirect effects can remain accessible through precision measurements.
The combined analysis of baryogenesis, electric dipole moment constraints, and indirect collider sensitivity reveals a finite and predictive region of parameter space where resonant enhancement, quasi-degeneracy, and CP violation simultaneously account for the observed baryon asymmetry while remaining consistent with current experimental bounds. The scaling behaviour of the asymmetry with the degeneracy parameter further quantifies the resonant nature of the hybrid mechanism.
Overall, hybrid resonant leptogenesis in renormalizable SO(10) unification provides a coherent link between the origin of neutrino masses, the baryon asymmetry of the Universe, and low-energy CP-violating observables. Future improvements in electric dipole moment searches, charged-lepton flavour violation experiments, and precision Higgs measurements will therefore play a crucial role in testing this framework.
@@ Acknowledgments
GG thanks UGC RUSA for carrying out this work.